# ReMoS: 3D Motion-Conditioned Reaction Synthesis for Two-Person Interactions

### v2

@@ 3.3 Losses and Training Details
This is the standard diffusion data term that minimizes the $\ell_{2}$-distance between the ground truth and the synthesized reactive motion, as $\mathcal{L}_{c}=\left\|X-X^{(0)}\right\|_{2}$. While the reconstruction loss provides the vital data term to drive the training, it does not enforce interaction synchronization between the actor and the reactor, or any kinematic constraints on the motion.
@@
We introduce the reaction loss to ensure accurate timing and spatial positioning of the reactor’s motion with respect to the actor’s motion. We calculate the ground-truth Euclidean distance $d(\cdot,\cdot)$ between each joint of the actor and the reactor at each frame and minimize the deviations from these distances for the actor and the synthesized reactor’s motions. We note that this distance term also implicitly constrains the reactor’s joints in frames that are less relevant for interactions, i.e., where the reactor’s joints are far from the actor’s joints. Further, the Euclidean distances for their hand joints in frames with no hand-based contact become drastically high. To mitigate these concerns, we use an exponentially decaying distance-aware weight: $\operatorname{exp}\left(-d(x_{n,j},y_{n,j})\right)$, at each frame $n$, to focus more on the reactor’s joints that are closer to the actor, thus more relevant for interaction (see Fig. 4). Thus, we get our reaction loss term as
$\mathcal{L}_{r}=\frac{1}{NJ}\sum_{n=1}^{N}\sum_{j=1}^{J}\operatorname{exp}\left(-d(x_{n,j},y_{n,j})\right)\cdot\\ \left|d(x_{n,j},y_{n,j})-d(x^{(0)}_{n,j},y_{n,j})\right|.$ (9)
@@
To maintain the kinematic plausibility of the generated motions, we follow existing literature on regularizers for bone length consistency, foot contacts, and temporal consistency [67, 55, 56]. To improve the temporal consistency of the motion [67], we minimize the joint velocities and joint accelerations between two consecutive frames of the ground-truth reactive motions, $X$, and the synthesized reactive motions, $\hat{X}$, defined as
$\mathcal{L}_{vel}=\frac{1}{N-1}\sum_{n=0}^{N-1}\bigl{(}(X^{n+1}-X^{n})-\\ (\hat{X}^{n+1}-\hat{X}^{n})\bigl{)}^{2},$ (10)
$\mathcal{L}_{acc}=\frac{1}{N-2}\sum_{n=0}^{N-2}\bigl{(}(X^{n+2}-2X^{n+1}+X^{n})-\\ (\hat{X}^{n+2}-2\hat{X}^{n+1}+\hat{X}^{n})\bigr{)}^{2},$ (11)
where $N$ is the total number of frames.
Additionally, we introduce a bone length loss consistency loss, $\mathcal{L}_{bone}$, to ensure that the synthesized reactor joint positions satisfy the skeleton consistency [42]. We define this loss as
$\mathcal{L}_{bone}=\left\|\mathbf{B}(X)-\mathbf{B}(\hat{X})\right\|^{2}_{2},$ (12)
Further, foot skating is a common artifact in motion synthesis [55, 56]. We constrain this by ensuring that the toe joint in contact with the ground plane has zero velocity. We use a binary foot contact loss [67, 66] on the foot joints of the synthesized pose to ensure that the output motion does not slide across the ground plane, defined as
$\mathcal{L}_{foot}=\frac{1}{N-1}\sum_{n=0}^{N-1}\left\|\left(\hat{X}^{n+1}-\hat{X}^{n}\right)\cdot\hat{\mathbbm{1}}^{n}_{foot}\right\|^{2}_{2},$ (13)
where $\hat{\mathbbm{1}}^{n}_{foot}\in\left\{0,1\right\}$ is the foot-ground contact indicator for the synthesized reactive motion $\hat{X}^{n}$ at each frame $n$. Our kinematic loss term is a weighted sum of the joint velocity loss $\mathcal{L}_{vel}$, the joint acceleration loss $\mathcal{L}_{acc}$, the bone length consistency loss $\mathcal{L}_{bone}$, and the foot skating loss $\mathcal{L}_{foot}$, as
@@
where $\lambda_{v}$, $\lambda_{a}$, $\lambda_{b}$ and $\lambda_{f}$ are scalar weights.
@@ 3.4 Inference Time Spatial Guidance
While our trained method synthesizes plausible reactive motions to diverse actor sequences without the need for additional annotations, we notice that for fast-paced contact-heavy interactions, the reactor’s body often misaligns spatially with the actor, especially at the arm joints. This, in turn, affects the synthesis of the finger joints in the second stage of our cascaded diffusion. Diffusion models give us the ability to guide the inference process using guidance functions that can provide the gradients to nudge the sampling process towards a certain direction [35, 51]. We use this strategy to guide our synthesized motion further to improve the spatial alignment with the actor. We design a guidance function $G_{A}\left(\phi,\hat{\phi}\right)\in\mathbb{R}^{N\times J_{A}\times 3}$, which minimizes the distance between the $J_{A}$ arm joints of the actor ($\phi$) and the $J_{A}$ arm joints of the synthesized reactor ($\hat{\phi}$). Specifically, we re-apply our interaction mask $\mathbbm{1}_{H}$ to determine which of the reactor’s hands are more likely to be interacting with the actor, and minimize the distances between the arm joints of the corresponding sides as $G_{A}=\operatorname*{arg\,min}_{\hat{\phi}}\left(\mathbbm{1}_{H}\left\|\phi-\hat{\phi}\right\|\right)$. We then plug $G_{A}$ into the denoising pipeline of our body diffusion module, as
$X_{B}^{(0)}=X_{B}^{(0)}-\gamma\nabla_{X_{B}^{(0)}}G_{A}\left(\phi,\hat{\phi}\right),$ (15)
where $\gamma$ is an emperically chosen guidance scale.
@@ 4 ReMoCap Dataset
To facilitate research on contact-based two-person interactions with finger-level articulations, we propose the ReMoCap dataset. This dataset captures complex two-person interactions in two different scenarios: the fast-paced, swing-style Lindy Hop dancing, and the martial arts of Ninjutsu. These two types of motions are well-suitable for studying physical interactions due to the following reasons. First, the spectrum of two-person motions in the aforementioned scenarios is significantly diverse in terms of intensity, speed and style. Second, choreographed motion generation is developing as an active field of research and we envision this dataset to facilitate research in interactive applications such as performing in virtual reality [9] and remote tutoring [52].
@@
We invited $4$ trained Lindy Hop dancers and $5$ trained Ninjutsu artists to perform diverse interactive movements in a multi-view capture studio. We tracked their skeletons using a commercially available markerless, multi-view motion-capture system [8], which tracks $93$ degrees-of-freedom driving $69$ body joints. Capturing motions in a markerless setting enables the performers to move uninhibited, while also making the data suitable for training monocular or multi-view motion capture methods under severe inter-person occlusions. We also capture RGB videos from $120$ camera views for each sequence. The dataset includes 3D skeleton poses with full-body and finger annotations, foreground-background segmentation masks, and 3D surface reconstructions of the subjects. We capture sequences of multiple lengths totaling ${\sim}275.7\mathrm{K}$ frames ($2.04$h) of motion data from each view at a frame rate of $50$ fps for the Lindy Hop and $25$ fps for the Ninjutsu. Out of all the frames, around $150\mathrm{K}$ frames have hand-based interactions between the two characters where the closest distance between the finger joints of the actor and the reactor is within $50$ mm (more statistics in the appendix).
@@
Table 1 shows a comprehensive comparison of ReMoCap with existing multi-person interaction datasets. Existing datasets consisting of two-person interactions such as SBU [76], K3HI [32], NTU-26 [43] and 2C [54] are limited in size and motion capture quality. They typically feature simple actions (such as handshakes, punching, pushing and kicking) with weak interactions (without hand motion capture). The recent ExPI [24] dataset captures Lindy Hop aerial sequences to model more complex interactions, and the InterHuman [42] dataset features both daily motions (e.g., passing objects, greeting, communicating) and professional motions (e.g., Taekwondo, Latin dance, boxing). However, these datasets do not provide finger-level motion capture data, i.e. a key requirement to intricately model inter-human performances. Two-person dance datasets, such as DuetDance [39] and DanceDB [52], also lack the hand motion data needed for modeling interactions. Only the concurrent work Duolando [59] (not published at submission time) proposes the DD100 dataset with strong interactions between two dancers and provides hand motion data.
@@ 5 Experiments and Results
We conduct comprehensive experiments to evaluate ReMoS on multiple two-person datasets covering a wide range of interaction scenarios. We perform training and evaluation on multiple large-scale datasets, including our proposed ReMoCap, 2C [54], and ExPI [24]. We provide the implementation details, report quantitative and qualitative comparisons on standard evaluation metrics and through a user study, report ablations, and show how to apply ReMoS as a motion editing tool. Details on the dataset preparation are in the appendix.
@@
5.1 Implementation Details
As a pre-processing step, we normalize the two-person body poses by translating the actor’s motion $Y$ and the reactor’s motion $X$ together, such that the root joint of $Y$ is at the global origin for all $N$ frames. We then compute the relative 3D joint coordinates of $X_{B}$ and $Y_{B}$ w.r.t. the root joint of $Y_{B}$. For the individual hands, we compute the relative 3D joint coordinates of $X_{H}$ and $Y_{H}$ w.r.t. the corresponding wrist joints of $X_{H}$ and $Y_{H}$. Using normalized inter-person root relative joint positions benefits the stability and convergence of our model. We train ReMoS for about $64\mathrm{K}$ iterations on both sequences of ReMoCap using the Adam optimizer [36], with a base learning rate of $10^{-5}$ and a batch size of $64$. We decay the learning rate using a Step LR scheduler with a step size of $5$ epochs and a decay rate of $0.99$. We use $d=256$ for our latent embedding representation and use $6$ layers in our transformer decoder with $4$ heads for calculating the attention. We train ReMoS with diffusion steps $T$ set to $500$, and the rate of diffusion, $\bar{\alpha}$, changing linearly from $0.0002$ to $0.02$. The training takes around $8$ and $11$ hours on an NVIDIA RTX A4000 GPU for Lindy Hop and Ninjutsu, respectively. We set $\lambda_{c},\lambda_{r},\lambda_{v}=10.0$, $\lambda_{a},\lambda_{k},\lambda_{b}=1.0$ and $\lambda_{f}=20.0$ (only applied after $100$ training epochs) as the loss term weights in Eqns. 8 and 14. We set the guidance scale $\gamma=10^{-3}$ in Eqn. 15. Similar training details on the other datasets are available in the appendix.
5.2 Baselines and Ablated Versions
For baselines, we choose the closest motion synthesis methods in a two-person setting, namely, MixNMatch [20], InterFormer [11], ComMDM [53], Role-Aware Interaction Generation (RAIG) [64], and InterGen [42]. InterFormer [11] was originally trained in a reactive motion synthesis setting without additional annotations, such as action labels or text descriptions, and we maintain this setup. For the other methods, we mask out their input text/label embeddings to comply with our annotation-free setting. We re-train these methods on ReMoCap with a thorough hyper-parameter search and report their best performances. More training details are in the appendix.
@@ 5.2 Baselines and Ablated Versions
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    W/o reaction loss. Removing the reaction loss (Eqn. 9) in training.
  • $\bullet$
    W/o spatial guidance. Removing spatial guidance $G_{A}$ (Sec 3.4) in inference.
5.3 Quantitative Evaluation
We evaluate ReMoS on the standard evaluation metrics. To measure the mean joint-level performance, we report the mean per-joint positional error (MPJPE) in mm, and the mean per-frame, per-joint velocity error (MPJVE) in mm on the synthesized motion by measuring the temporal consistency deviation from the ground truth 3D pose following [56]. We measure the Fréchet Inception Distance (FID) [27] to compare the distribution gap between the generated motion embedding spaces of the different methods and the ground truth motion embedding space. As part of our method focuses on generating hand motions, we measure the FID score of body and hands separately for ReMoCap. We also compute the latent variance of the generated motions (Diversity) [66, 79]. Table 2 reports the quantitative evaluation of ReMoS with its baselines and ablations on ReMoCap. ReMoS achieves state-of-the-art performance for both the Lindy Hop pair-dance setting and the Ninjutsu setting. We observe that methods using denoising diffusion have an overall higher Diversity compared to transformer-based methods [11], or GAN-based methods [20]. This enforces the variability claims of denoising diffusion models. We notice atleast $20\%$ improvement in the MPJPE and around $40\%$ improvement in the FID score for the reactor’s body motion when using the proposed CoST-XA mechanism, thus confirming that this mechanism helps in learning the fine-grained, inter-person dependencies across time. We also notice an improvement of almost $50\%$ in the FID scores of the reactor’s hand motion when using the cascaded strategy and the reaction loss. Further, the spatial guidance function fine-tunes the overall reaction and improves MPJPE and MPJVE. Note that the MPJPE values are overall higher in the Ninjutsu setting as compared to the Lindy Hop since the trajectory of the reactor is more diverse for Ninjutsu. We also report the evaluation of ReMoS with its baselines on the ExPI [24] and the 2C [54] datasets in Table 3. These datasets do not provide hand motions, so we only evaluate the reactor’s body motions. ReMoS achieves state-of-the-art performance in aforementioned metrics in these datasets.
Figure 5: Qualitative Results and Applications. Top: Visual comparison of ReMoS with GT and baselines. The reactor (blue) synthesized by ReMoS has the most plausible alignment with the actor (red) compared to our baselines. Bottom: Using ReMoS as an interactive motion editing tool. Bottom left: We show a pose-completion application where we manually control the right-hand wrist joint of the reactor and let ReMoS synthesize the remaining body joints conditioned on the actor. Bottom right: ReMoS used to synthesize reactor’s motion in-between start and end frames.
@@ 5.4 User Study
We evaluate the visual quality of our captured ground truth data and the generated reactive motions through a user study. We show participants $26$ different interaction sequences across the ground truth, our method, and its baselines. For each sequence, we randomly show them three methods side-by-side and ask them to rate the 3D motions they observe in terms of (a) the reactor’s motion quality, irrespective of the actor’s motion (Motion Quality), and (b) the plausibility of the reaction given the actor’s motion (Reaction Plausibility). We ask the participants to rank these motions from ‘1’ (lowest score) to ‘5’ (best score) in a five-point Likert scale. Table 4 reports the mean scores from the responses of $40$ participants, excluding the responses that did not pass our validation checks. We notice the ground truth motions in ReMoCap achieve the highest score of around $96\%$, indicating that participants perceive our captured ground truth motions to be natural-looking, with realistic interactions. ReMoS has the second best ranking of around $78\%$, which is almost $21\%$ higher than the baselines, showing that it is preferred more over the baselines.

### v3

@@ 3.3 Losses and Training Details
This is the standard diffusion data term that minimizes the $\ell_{2}$-distance between the ground truth and the synthesized reactive motion, as $\mathcal{L}_{c}=\left\|X-X^{\left(0\right)}\right\|_{2}$. While the reconstruction loss provides the vital data term to drive the training, it does not enforce interaction synchronization between the actor and the reactor, or any kinematic constraints on the motion.
@@
We introduce the reaction loss to ensure accurate timing and spatial positioning of the reactor’s motion with respect to the actor’s motion. We calculate the ground-truth Euclidean distance $d\left(\cdot,\cdot\right)$ between each joint of the actor and the reactor at each frame and minimize the deviations from these distances for the actor and the synthesized reactor’s motions. We note that this distance term also implicitly constrains the reactor’s joints in frames that are less relevant for interactions, i.e., where the reactor’s joints are far from the actor’s joints. Further, the Euclidean distances for their hand joints in frames with no hand-based contact become drastically high. To mitigate these concerns, we use an exponentially decaying distance-aware weight $\operatorname{exp}\left(-d\left(x_{n,j},y_{n,j}\right)\right)$ at each frame $n$ to focus more on the reactor’s joints that are closer to the actor, and therefore more relevant for interaction (see Fig. 4). Thus, we get our reaction loss term as
$\mathcal{L}_{r}=\frac{1}{NJ}\sum_{n=1}^{N}\sum_{j=1}^{J}\operatorname{exp}\left(-d\left(x_{n,j},y_{n,j}\right)\right)\cdot\left|d\left(x_{n,j},y_{n,j}\right)-d\left(x^{\left(0\right)}_{n,j},y_{n,j}\right)\right|.$ (9)
@@
To maintain the kinematic plausibility of the generated motions, we follow existing literature on regularizers for bone length consistency, foot contacts, and temporal consistency [68, 56, 57]. Our kinematic loss term is a weighted sum of the joint velocity loss $\mathcal{L}_{vel}$, the joint acceleration loss $\mathcal{L}_{acc}$, the bone length consistency loss $\mathcal{L}_{bone}$, and the foot sliding loss $\mathcal{L}_{foot}$, as
@@
where $\lambda_{v}$, $\lambda_{a}$, $\lambda_{b}$ and $\lambda_{f}$ are scalar weights (more details in the appendix).
@@ 3.4 Inference Time Spatial Guidance
While our method synthesizes plausible reactive motions, it can sometimes spatially misalign the reactor’s body to the actor’s, espeically at the arm joints, for fast-paced, contact-heavy interactions. This, in turn, affects finger-joint synthesis in the second stage of our cascaded diffusion. To improve spatial alignment, we leverage guidance functions [35, 52] that can provide gradients to nudge the sampling process towards a certain direction. We design a guidance function $G\left(\phi,\hat{\phi}\right)\in\mathbb{R}^{N\times J_{A}\times 3}$, which minimizes the distance between the $J_{A}$ arm joints of the actor ($\phi$) and the $J_{A}$ arm joints of the synthesized reactor ($\hat{\phi}$). Specifically, we re-apply our interaction masks $\mathbbm{1}_{H_{A}}$ and $\mathbbm{1}_{H_{R}}$ to determine which of the reactor’s hands are more likely to be interacting with the actor, and minimize the distances between the arm joints of the corresponding sides as $G=\operatorname*{arg\,min}_{\hat{\phi}}\left(\left\|\mathbbm{1}_{H_{A}}\odot\phi-\mathbbm{1}_{H_{R}}\odot\hat{\phi}\right\|\right)$. We then plug $G$ into the denoising pipeline of our body diffusion module, as
$X_{B}^{\left(0\right)}=X_{B}^{\left(0\right)}-\gamma\nabla_{X_{B}^{\left(0\right)}}G\left(\phi,\hat{\phi}\right),\qquad\gamma=\textrm{ guidance scale}.$ (11)
@@ 4 ReMoCap Dataset
We propose the ReMoCap dataset to facilitate research on contact-based two-person interactions with finger-level articulations. It contains complex two-person interactions in two separate scenarios: the fast-paced, swing-style Lindy Hop dancing, and the martial arts of Ninjutsu. We consider these two types of motions for studying physical interactions for two main reasons. First, these motions encompass a significantly diverse spectrum of motions in terms of intensity, speed, and style. Second, choreographed motion generation is developing as an active field of research and we envision this dataset to facilitate research in interactive applications such as performing in virtual reality [9] and remote tutoring [53].
@@
We invited $4$ trained Lindy Hop dancers and $5$ trained Ninjutsu artists to perform diverse interactions in a multi-view capture studio. We tracked their motions using a commercially available markerless, multi-view motion-capture system [8], which tracks $93$ degrees-of-freedom driving $69$ body joints. Capturing motions in a markerless setting enables the performers to move uninhibited, while also making the data suitable for training monocular or multi-view motion capture methods under severe inter-person occlusions. We also capture RGB videos from $120$ camera views for each sequence. The dataset includes 3D skeleton poses with full-body and finger annotations, foreground-background segmentation masks, and 3D surface reconstructions of the subjects. We capture sequences of multiple lengths totaling ${\sim}275.7\mathrm{K}$ frames ($2.04$ hours) of motion data from each view at a frame rate of $50$ fps for the Lindy Hop and $25$ fps for the Ninjutsu. Out of all the frames, around $150\mathrm{K}$ frames have hand-based interactions between the two characters where the closest distance between the finger joints of the actor and the reactor is within $50$ mm. We present more dataset statistics in the appendix.
@@
Table 1 shows a comprehensive comparison of ReMoCap with existing multi-person interaction datasets. Existing datasets consisting of two-person interactions, such as SBU [78], K3HI [32], NTU-26 [43] and 2C [55], are limited in size and motion capture quality. They typically feature simple actions, such as handshakes, punching, pushing and kicking, with weak interactions, and do not capture hand motions. The recent ExPI [24] dataset captures Lindy Hop aerial sequences to model more complex interactions, and the InterHuman [42] dataset features both daily motions (e.g., passing objects, greeting, communicating) and professional motions (e.g., Taekwondo, Latin dance, boxing). However, these datasets do not provide finger-level motion capture data, which is a key requirement to intricately model inter-human activities. Two-person dance datasets, such as DuetDance [39] and DanceDB [53], also lack the hand motion data needed for modeling interactions. Only the concurrent work, Duolando [60]111unpublished at the time of our paper submission proposes the DD100 dataset with strong interactions between two dancers and provides hand motion data.
@@ 5 Experiments and Results
We conduct comprehensive experiments to evaluate ReMoS on multiple two-person datasets covering a wide range of interaction scenarios. We perform training and evaluation on multiple large-scale datasets, including 2C [55], ExPI [24], and our proposed ReMoCap. We provide the implementation details, report quantitative and qualitative comparisons on standard evaluation metrics and through a user study, report ablations, and show how to apply ReMoS as a motion editing tool. We provide details on the dataset preparation in the appendix.
5.1 Implementation Details
As a pre-processing step, we normalize the two-person body poses by translating the actor’s motion $Y$ and the reactor’s motion $X$ together, such that the root joint of $Y$ is at the global origin for all $N$ frames. We then compute the relative 3D joint coordinates of $X_{B}$ and $Y_{B}$ w.r.t. the root joint of $Y_{B}$. For each hand, we compute the relative 3D joint coordinates of $X_{H}$ and $Y_{H}$ w.r.t. the corresponding wrist joints of $X_{H}$ and $Y_{H}$. Using normalized inter-person, root-relative joint positions benefits the stability and convergence of our model. ReMoS uses $T=500$ diffusion steps where $\bar{\alpha}$ changes linearly from $0.0002$ to $0.02$. We train for about $64\mathrm{K}$ iterations on both sequences of ReMoCap using Adam [36] with a base learning rate of $10^{-5}$ and a batch size of $64$. We decay the learning rate using a Step LR scheduler with a step size of $5$ epochs and a decay rate of $0.99$. We use $d=256$ for our latent embedding representation and use $6$ layers in our transformer decoder with $4$ heads for calculating the attention. The training takes around $8$ and $11$ hours on an NVIDIA RTX A4000 GPU for Lindy Hop and Ninjutsu, respectively. The inference time is ${\sim}24s$ to generate $50$ frames of full-body and finger motions ($12.5s$ for body synthesis and then $11.5s$ for hand synthesis). We set $\lambda_{c},\lambda_{r},\lambda_{v}=10.0$, $\lambda_{a},\lambda_{k},\lambda_{b}=1.0$ and $\lambda_{f}=20.0$ (only applied after $100$ training epochs) as the loss term weights in Eqns. 8 and 10. We set the guidance scale $\gamma=10^{-3}$ in Eqn. 11.
5.2 Baselines and Ablated Versions
For baselines, we choose the closest motion synthesis methods in a two-person setting, namely, MixNMatch [20], InterFormer [11], ComMDM [54], Role-Aware Interaction Generation (RAIG) [65], and InterGen [42]. InterFormer [11] was originally trained in a reactive motion synthesis setting without additional annotations, such as action labels or text descriptions, and we maintain this setup. For the other methods, we mask out their input text/label embeddings to comply with our annotation-free setting. We re-train these methods on ReMoCap with a thorough hyper-parameter search and report their best performances. We provide more training details in the appendix.
@@ 5.2 Baselines and Ablated Versions
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    w/o reaction loss. Removing the reaction loss (Eqn. 9) in training.
  • $\bullet$
    w/o spatial guidance. Removing spatial guidance $G_{A}$ (Sec. 3.4) in inference.
5.3 Quantitative Evaluation
We evaluate ReMoS on standard evaluation metrics. We measure the temporal consistency deviation from the ground truth 3D pose following [57] to report the mean per-joint positional error (MPJPE) and the mean per-frame, per-joint velocity error (MPJVE), both in mm, on the synthesized motions. We measure the Fréchet Inception Distance (FID) [27] to compare the distribution gap between the embedding spaces of the generated and ground-truth motions. As part of our method focuses on generating hand motions, we measure the FID score of body and hands separately for ReMoCap. We also compute the latent variance of the generated motions (Diversity) [67, 81]. Table 2 reports the quantitative evaluation of ReMoS with its baselines and ablations on ReMoCap. ReMoS achieves state-of-the-art performance for both the Lindy Hop pair-dance setting and the Ninjutsu setting. We observe that methods using denoising diffusion have higher Diversity compared to transformer-based [11] or GAN-based [20] methods. This enforces the variability claims of denoising diffusion models. We note at least $20\%$ improvement in the MPJPE and around $40\%$ improvement in the FID score for the reactor’s body motion when using the proposed CoST-XA mechanism, confirming its benefit in learning fine-grained, inter-person dependencies across time. We also note an almost $50\%$ improvement in the FID scores of the reactor’s hand motion when using the cascaded strategy and the reaction loss. Further, the spatial guidance function fine-tuning improves MPJPE and MPJVE. The MPJPE values are overall higher in Ninjutsu than in Lindy Hop as the trajectory of the reactor is more diverse for Ninjutsu. We also report the evaluation of ReMoS with its baselines on the ExPI [24] and the 2C [55] datasets in Table 4. These datasets do not provide hand motions, so we only evaluate the reactor’s body motions, and report state-of-the-art performance of ReMoS.
@@ 5.4 User Study
We evaluate the visual quality of our captured ground truth data and the generated reactive motions through a user study. We show participants $26$ different interaction sequences across the ground truth, our method, and its baselines. For each sequence, we randomly show them three methods side-by-side and ask them to rate the 3D motions they observe in terms of (a) the reactor’s motion quality, irrespective of the actor’s motion (Motion Quality), and (b) the plausibility of the reaction given the actor’s motion (Reaction Plausibility). We ask the participants to rank these motions from ‘1’ (worst score) to ‘5’ (best score) in a five-point Likert scale. Table 4 reports the mean scores from the responses of $40$ participants, excluding the responses that did not pass our validation checks. We notice the ground truth motions in ReMoCap achieve the highest score of around $96\%$, indicating that participants perceive our captured ground truth motions to be natural-looking, with realistic interactions. ReMoS has the second best ranking of around $78\%$, which is almost $21\%$ higher than the baselines, showing that it is preferred more over the baselines.